6. Suppose the following data describe a nations Year 2 Year I 310 million 150 million 8 percent 315 million Population Labor Force Unemployment rate 160 million 7 percent How many people ure unemployed in each year? (5 pts) b. How many people are employed in each ycar? (5 ps) c. If in Year 1 the full-employment level of output was $3,000 billion and the full-employment rate of unemployment was 5.5%, how much potential GDP was lost in Year l based on Okuns Law? (4 pls) d. List and briely discuss the different types of uncmployment mentioned in class. (6 pts)
Answer #1

1- In year 1, No. of unemployed people=8% of 150 million=12 million

In year 2, no of unemployed people= 7% of 160 million=11.2 million

2- In year 1, Employed people=150-12=138 million

In year 2, employed people=160-11.2=148.8 million.

3- Okun's law states that for every 1% increase in the employment rate, gross domestic product increases 3%.

4- Types of unemployment-

a) Structural- It occurs because of absence of demand of certain type of worker in labor market. This happens when there is mismatch between skills workers have and skills employees want.

b) Frictional- Frictional unemployment occurs because of the normal turnover in the labor market and the time it takes for workers to find new jobs. Throughout the course of the year in the labor market, some workers change jobs. When they do, it takes time to match up potential employees with new employers. Even if there are enough workers to satisfy every job opening, it takes time for workers to learn about these new job opportunities, and for them to be considered, interviewed and hired.

c) Cyclical- Over time, the economy experiences many ups and downs. Cyclical unemployment occurs because of these ups and downs. When the economy enters a recession, many of the jobs lost are considered cyclical unemployment.